Grade 12Physical Chemistry

The relative decrease in vapour pressure of an aqueous solution containing 2 mole NaCl in 3 mole H2O is 0.40 . on reaction with agno3 this solution will form

Geetha
8 Years agoGrade 12

2 Answers

sid dwivedi
6 Years ago
From RLVP , we get
 
.4 = i(1)/ i(1) + 3
 
So i = 2 
 
Dissociation of  Nacl
 
Na^+  +   cl^-
 
i = 1 + alpha 
 
Then  1+ alpha = 2   so alpha = 1 
 
 
Means 100% ionization
 
 
 Means 100% ionization
Vikas TU
6 Years ago
Dear student 
Relative lowering in vapour pressure 
del p / p^0 = inb /(inb +na +...)
del p = p^0 -p
For Agcl = 3mol - 2mol = 1mole gcl 
Hope this helps 
Good Luck